OpenStudy (rootbeer003):

Which pair of numbers is not a factor pair of 40? 4 and 10 –4 and –10 –8 and 5 20 and 2

OpenStudy (rootbeer003):

AYEEEEEE

OpenStudy (anonymous):

oh no!!!

OpenStudy (rootbeer003):

WELCOME

OpenStudy (anonymous):

what is \(-8\times 5\)?

OpenStudy (rootbeer003):

-45

OpenStudy (anonymous):

hint, it is not \(40\)

OpenStudy (anonymous):

:D well for the third one, what does -8 x 5 make?

OpenStudy (rootbeer003):

40*

OpenStudy (frozen15):

@satellite73 I couldn't message you but are you good at g.um? English sorry for intrupting the question

OpenStudy (anonymous):

no not 40

OpenStudy (anonymous):

you got one minus sign right? where did it go?

OpenStudy (rootbeer003):

OHHHH IM DUMBB

OpenStudy (anonymous):

Yup

OpenStudy (rootbeer003):

WOOT

OpenStudy (rootbeer003):

um no satelite get back here

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

How is that wrong???

OpenStudy (rootbeer003):

Which of the following numbers is prime? 111 121 131 141

OpenStudy (anonymous):

\[-8\times 5=-40\neq 40\]

OpenStudy (anonymous):

oh another question

OpenStudy (rootbeer003):

im not done

OpenStudy (anonymous):

111 is divisible by 3 since \(1+1+1=3\)

OpenStudy (anonymous):

\[121=11^2\] not prime

OpenStudy (rootbeer003):

but it asks which is prime

OpenStudy (anonymous):

141 is also divisible by 3 since \(1+4+1=6\)

OpenStudy (anonymous):

as usual, it is C 131

OpenStudy (anonymous):

you now how to tell if a number is divisible by 3?

OpenStudy (rootbeer003):

Determine the prime factorization of 99. 34 34• 2 • 5 34• 11 32• 11

OpenStudy (anonymous):

??

OpenStudy (rootbeer003):

HOW

OpenStudy (anonymous):

\[99=9\times 11\] right ?

OpenStudy (rootbeer003):

ye

OpenStudy (anonymous):

and since \(9=3\times 3\) that means \[99=3\times 3\times 11=3^2\times 11\] which does not see bot be one of your choices

OpenStudy (rootbeer003):

mhm

OpenStudy (anonymous):

unless your last choice is \(3^2\times 11\) and not \(32\times 11\) which is what you wrote

OpenStudy (rootbeer003):

mhmm

OpenStudy (rootbeer003):

so d

OpenStudy (anonymous):

yeah D

OpenStudy (rootbeer003):

alright

OpenStudy (rootbeer003):

Determine the prime factorization of 476. 22 • 32 • 13 22 • 7 • 19 22• 7 • 17 2 • 35

OpenStudy (anonymous):

donkey work

OpenStudy (rootbeer003):

hay im tryin

OpenStudy (anonymous):

want to do it the long boring way ?

OpenStudy (rootbeer003):

no

OpenStudy (rootbeer003):

wanna do the ' i learn ' it way

OpenStudy (rootbeer003):

>:)

OpenStudy (anonymous):

ok then it is clear that \(476\) is divisble by 2 right, because it is even

OpenStudy (rootbeer003):

yass

OpenStudy (anonymous):

ok so step one is find half of \(476\) i.e. divide it by 2 what do you get?

OpenStudy (rootbeer003):

238

OpenStudy (anonymous):

ok so step one \[476=2\times 238\] but \(238\) is also even, what is half of 238?

OpenStudy (rootbeer003):

119

OpenStudy (anonymous):

ok so we are at \[476=2^2\times 119\] 119 is not even, s it is not divisible by 2

OpenStudy (anonymous):

\[1+1+9=11\] so it is not divisible by 3

OpenStudy (anonymous):

it is obviously not divisible by 5

OpenStudy (anonymous):

try 7

OpenStudy (rootbeer003):

k

OpenStudy (rootbeer003):

17!

OpenStudy (rootbeer003):

took me awhile to process that

OpenStudy (anonymous):

ok so \[476=2^2\times 7\times 17\]

OpenStudy (rootbeer003):

cccccc

OpenStudy (rootbeer003):

hello?

OpenStudy (anonymous):

hi!

OpenStudy (rootbeer003):

Find the greatest common factor of the pair of numbers. 104 and 260 4 13 20 52

OpenStudy (anonymous):

yeah C

OpenStudy (anonymous):

\[104=2\times 52\] right? does \(52\) go in to 260 evenly?

OpenStudy (rootbeer003):

ya

OpenStudy (anonymous):

then pick that one

OpenStudy (rootbeer003):

dddddddddddddd

OpenStudy (rootbeer003):

?

OpenStudy (anonymous):

do i have to do everything? yes D

OpenStudy (rootbeer003):

lagg

OpenStudy (rootbeer003):

Find the greatest common factor of the pair of numbers. 348 and 426 12 6 9 3

OpenStudy (rootbeer003):

ill look into this one

OpenStudy (anonymous):

it has be be even right, because both numbers are even

OpenStudy (anonymous):

see if it is 12, if not, it is 6

OpenStudy (rootbeer003):

well i dont want to get it wrong

OpenStudy (rootbeer003):

um 6

OpenStudy (rootbeer003):

cause its even

OpenStudy (anonymous):

lage real bad hold on

OpenStudy (anonymous):

you are right, 6

OpenStudy (rootbeer003):

ready 4 some more?

OpenStudy (anonymous):

if i can get past this lage

OpenStudy (rootbeer003):

simplify

OpenStudy (anonymous):

\[\frac{a^9}{a^4}=a^{9-4}\]

OpenStudy (rootbeer003):

c

OpenStudy (rootbeer003):

waitt

OpenStudy (rootbeer003):

a

OpenStudy (anonymous):

A

OpenStudy (rootbeer003):

iz a?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i have never heard of this but the only answer that makes sense is "power of products"

OpenStudy (rootbeer003):

huh k

OpenStudy (rootbeer003):